OpenStudy (anonymous):

Common volume for z=3, sphere, x^2+y^2=3*x,cylinder. Use trippel integral and spherical coordinates

OpenStudy (anonymous):

x^2-3x+___+y^2=0 x^2-3x+9/4+y^2=9/4 (x-3/2)^2+y^2=9/4 in the xy plane you have a cylinder center at (3/2,0) with radius 3/2 thats going to go up to 3 on the z-axis i dont know if i am doing this correctly but i dont see a sphere but i can tell you \[\int\limits_{0}^{\pi/2}\int\limits_{0}^{2\pi}\int\limits_{0}^{3/2}r^2\sin(\Theta)drd (\Theta)d (\phi)\]

OpenStudy (anonymous):

i am thinking you need to multiply the intergal by 2 to get the full sphere and not just the top cap, but if its a sphere and z=3 than the spheres radius =3 IMO so on that r=3/2 it could be 3 and the PHI should go from 0topi and then you woulnt have to multiply it by 2.

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

are you doing this online or in a book?

OpenStudy (anonymous):

a book

OpenStudy (anonymous):

ok the only other thing i could think of is (x-3/2)^2+y^2=z^2/4 that will give you the drawing of a phere

OpenStudy (anonymous):

have cal. with cylinder coordinates with correct answer 18pi-24

OpenStudy (anonymous):

cylindrical coordinated dzrdrd(theta) is the easy way to do it and spherical coor is just another way dz=0 to 3 r=0 to 3/2 and d(theta)=0 to 2pi right? thats what you did for this one

OpenStudy (anonymous):

no. theta from 0 to pi/2. r from 0 to 3*cos(theta) and z from 0 to sqrt(9-r^2) . result times 4

OpenStudy (anonymous):

ok now it makes sense but when you said x^2+y^2=3*x im just curious where your getting the 9-r^2 and x^2+y^2=r^2 maybe i didnt read the question in full

OpenStudy (anonymous):

phere: x^2+y^2+z^2=3^2 and x^2+y^2=r^2 here you find z

OpenStudy (anonymous):

did you mean 3^2 not 3*x in the original eq

OpenStudy (anonymous):

no.sphere has r=3

OpenStudy (anonymous):

OOO yes than that would have been easier to do, sorry this question took so long, it would have been easier if i would have noticed that in the begining sorry

OpenStudy (anonymous):

Ok. No problems, but i wonder about this spherical coord. My problem is to find the correct phi

OpenStudy (anonymous):

well i think i can help. so phi starts at xy=0 or on the z axis(positive) and moves towards the xy axis until z=0 and the phi there is pi/2. now thats moving 90 degrees. if you move 180 degrees from the z-axis to the -z-axis the phi is pi.

OpenStudy (anonymous):

i found the easiet way to do spereical coords is to to the r first usually going to be something involving (theta) then do the Dtheta next to get rid of the annoying sin and cos and in the end your left with phi which is really easy to plug in

OpenStudy (anonymous):

I have found 2 phi's, but when I use them I got 18*pi-27 or 18*pi-36 wrong answer

OpenStudy (anonymous):

what were your two phis

OpenStudy (anonymous):

sorry. it was cola time. One phi was pi/2-theta. number two very complicated arccos angle